OpenStudy (anonymous):

solve...0.7x+5le1.5x-2

OpenStudy (anonymous):

\[0.7x+5\le1.5x-2\]

OpenStudy (a_clan):

simple question. you just have to add and multiply terms

OpenStudy (anonymous):

1.2\[\ge\]1.3

OpenStudy (anonymous):

that didnt work to well

OpenStudy (a_clan):

add 2 both sides subtract 0.7x both sides

OpenStudy (a_clan):

your objective should be to bring x, alone, on one side of the inequality

OpenStudy (anonymous):

0.8 and 2.2

OpenStudy (a_clan):

0.7x + 7 <= 1.5x

OpenStudy (a_clan):

7 <= 1.5x - 0.7x

OpenStudy (a_clan):

7 <= 0.8x

OpenStudy (a_clan):

7/0.8 <= x

OpenStudy (a_clan):

or 70/8 <= x

OpenStudy (anonymous):

so it would be \[\left\{ x \left| \le70/8 \right| \right\}\]

OpenStudy (anonymous):

it says type and integer or a decimal

OpenStudy (anonymous):

an integer or a decimal

OpenStudy (anonymous):

\[\text{Reduce}\left[\frac{7}{10}x+5\text{$<$=}\frac{3}{2}x-2\right]\to \left(x\geq \frac{35}{4}\right) \]Reduce is a Mathematica function.